# Contrastive Learning for Task-Independent SpeechLLM-Pretraining

### v1

@@ 5.4 Additional pretraining data
As before, we use the 10% finetuning subset to simulate the low-resource scenario. In preliminary experiments with lightweight contrastive pretraining on the embedding layer (Appendix H), we find that increasing the pretraining data leads to much larger improvements in the normalized average with contr-cos loss ($+33$) than with contr-wasser loss ($+6.10$). Therefore, we use contr-cos for subsequent experiments, where we perform contrastive learning on all layers with and without ASR loss, as these settings yielded the best results in the previous sections. As shown in Table 4, both models contr-cos-all+giga and contr-cos-all+asr+giga achieve an overall improvement of more than $20$ points compared to the same models trained with less pretraining data, reaching $97.43$ and $102.3$ points respectively. Remarkably, despite being finetuned on only 10% of the task-specific data, these models surpass both specialized models, the Seamless model (Communication et al., 2023) for ST and the You et al. (2022) model for SQA. Moreover, they outperform the ASR-pretraining baseline even on the ASR task.
@@
We also evaluate our best-performing models in a higher resource setting. We finetune contr-cos-all+giga and contr-cos-all+asr+giga on 100% of the task-specific data. As expected, this leads to even stronger results. But while the ASR loss provides significant improvements in the low-resource settings, when finetuning with 100% of the available data, the contrastive loss alone performs on par with the combined approach. We hypothesize that the ASR loss is more sensitive to the out-of-domain Giga pretraining data than the contrastive loss, leading to no improvement on the ASR task.
We compare these models against blsp_lsm_7b (Wang et al., 2024a), which has a similar architecture to ours in that it trains only the projector while keeping other components frozen, and Qwen2-Audio-7b-Instruct (Chu et al., 2024), a strong baseline SpeechLLM. Notably, BLSP is finetuned on Must-C translation data, whereas for all other tasks, BLSP and Qwen2-Audio results reflect zero-shot performance. Across all tasks, our model outperforms both of these SpeechLLMs.
@@ 5.5 Impact of text-speech alignment on capturing paralinguistic features
Speech data is much richer than a sequence of words, and the pretrained speech encoder (HuBERT) might capture additional paralinguistic features. By aligning the richer speech embeddings more closely with text embeddings, the pretraining methods we explore could potentially train the projector to discard these additional features. To assess whether this occurs, we finetune our models on paralinguistic classification tasks before and after pretraining. In addition, we use the speech encoder with a linear classification head as a baseline.
The results demonstrate that models with contrastive and ASR-based pretraining perform comparably to the plain HuBERT model in predicting paralinguistic annotations, indicating that alignment-focused pretraining does not significantly affect the retention of paralinguistic information. Detailed accuracy results are presented in Table 12 in Appendix I. In addition, Table 13 shows that incorporating these features as auxiliary finetuning tasks does not notably alter model performance.
@@ 7 Limitations
We analyze whether contrastively pretrained models, designed to align speech and text, retain their ability to capture meta-speech information such as speaker characteristics or noise. While we find that these models do not lose this capacity, extending this analysis to generation tasks where such meta-speech features influence the output would provide a deeper understanding of their impact.
@@
Potential Risks
Enabling Language Models to understand speech makes them more easily and widely accessible. However, this also lowers the barrier for the potential misuse of these models. In addition, SpeechLLMs can reproduce biases seen in the training data. We try to minimize this risk by not training the LLM, so any safeguard measures trained into the LLM also apply to the SpeechLLM. However, the model could still be affected by adversarial attacks.
@@ Acknowledgments
Appendix A Mixed Next Word Prediction
For alignment between text and speech, we utilize the forced alignment algorithm described by Kurzinger et al. (2020), which leverages wav2vec 2.0 (Baevski et al., 2020).
Appendix B Hyperparameters and Model Details
@@
Q-Former Num Attention Heads 13
@@ Appendix C Data
For ASR and ST, we use the MustC v1 dataset (Di Gangi et al., 2019). We use the en-de portion as English ASR data, and we use language pairs en-de, en-fr, en-it, and en-es for ST. We exclude examples longer than 45 seconds for training and use the tst-common and tst-he test sets for evaluation.
@@
Contrastive pretraining aligns text and speech, which could potentially lead to the loss of paralinguistic features in the embeddings. Therefore, we also train and test our models on mls-eng-speaker-descriptions (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024). We use the mls-eng-speaker-descriptions train and test set. The available training data is quite large, so we subsample it to make training feasible. Since the test set is predefined, we sample the training data such that the test data size is 5% of the data. We let the model predict the following features: Speaking rate, gender, noise, reverberation, speech monotony, SDR noise, and PESQ speech quality.
Details on prompts can be found in Appendix D.
Appendix D Prompts
@@
(a) Loss during contr-cos-petraining
(b) Loss during contr-wasser-petraining
@@
Appendix E Contrastive Loss over Multiple Layers
This section analyses contrastive pretraining over multiple layers. Fig. 3 shows the contrastive loss in the pretraining phase. The later the layer, the lower the contrastive loss, with an exception of the embedding layer. For cosine contrastive training, these differences are significanlty bigger than for wasserstein contrastive training. Fig. 4 shows the finetuning performance on 10% on task-specific data after pretraining on different layers for the tasks ASR, ST and SQA. Here the losses for pretraining seem to correlated mostly with the performance of the model after finetuning: Models perform better after contrastive pretraining at later layers. However, for ASR, this trend is not so noticeable.
@@ Appendix E Contrastive Loss over Multiple Layers
10% no pretrain 23.78 19.869 69.722 31.345 41.859 -57.401
ASR pretrain 12.21 24.822 75.702 49.478 63.358 39.934
10% contr-cos-emb + mixed 14.26 25.96 76.605 51.968 68.395 46.614
contr-wasser-emb + mixed 14.7 24.145 75.045 48.012 61.48 26.531
mixed-contr-cos-last 19.02 23.023 73.92 46.285 58.557 3.859
mixed-contr-wasser-last 16.03 23.09 74.14 43.835 58.201 12.947
contr-cos-all + mixed 17.63 22.901 73.828 42.149 55.117 2.136
10% mixed-nwp with punctuation 15.76 24.229 75.085 47.651 61.36 23.586
mixed-nwp without punctuation 15.27 25.224 76.05 53.574 67.678 39.667
mixed nwp without punctuation + giga 14.91 26.071 77.172 56.988 70.935 51.673
Table 9: Ablation studies for models with mixed speech and text input. Metrics with * contribute to the normalized average as in Eq. 3.
@@ Appendix F Mixed Speech-Text Inputs
We test contrastive pretraining with mixed speech-text input (contr-*+mixed) and find that aligning the representation on the embedding layer works best (however, contrastive learning without mixed input still works better). We also experiment with mixed speech-text next word prediction (nwp). Here, removing punctuation seems to help the model understand the NWP task better. Finally, we also combine the NWP task with calculating the contrastive loss and summthe two losses, which also does not seem to help the model to align text and speech embeddings.
@@
10% ASR pretrain 12.21 24.822 75.704 49.478 63.358 39.934
10% contr-cos-emb 13.98 25.474 76.322 55.1 68.719 46.339
   + asr loss 13.18 25.773 76.892 56.124 69.229 52.467
   + mixed + nwp 14.68 25.282 76.175 48.855 63.922 36.406
contr-wasser-emb 15.99 26.096 77.085 55.04 68.166 44.024
   + asr loss 13.19 26.552 77.4 59.016 70.79 57.544
   + mixed + nwp 16 23.98 74.593 45 59.113 16.868
10% contr-cos-all 13.06 27.188 78.285 60.482 73.9 67.388
   + asr loss 11.68 27.608 78.94 64.157 76.396 78.577
   + mixed + nwp 22.08 23.16 74.195 48.474 62.126 2.073
contr-wasser-all 12.92 29.072 80.522 64.036 77.258 85.017
   + asr loss 17.63 29.064 80.26 66.727 77.565 70.784
   + mixed + nwp 15.72 24.13 74.993 51.205 64.254 27.515
@@ Appendix H Adding more pretraining data
N/A Specialized 06.54 30.99 80.02 64.1 77.1 N/A N/A 100
N/A BLSP-lslm-7b 44.51 28.70 78.68 5.60 21.82 N/A N/A -96.72
@@
   + giga 10.94 29.95 81.21 65.58 79.32 1.06 0.71 097.43
contr-cos-all + asr 11.68 27.61 78.94 64.16 76.40 1.08 0.91 078.58
   + giga 11.12 29.90 81.29 72.43 82.63 0.98 0.65 102.30
@@
Speech data is much richer than a sequence of words, and the pretrained speech encoder (HuBERT) might capture additional paralinguistic features. However, contrastive and ASR pretraining could potentially lead to discarding these paralinguistic features, when pulling speech embeddings closer to text embeddings. To assess this, we finetune our models on paralinguistic classification tasks before and after pretraining. We also compare our models against a HuBERT + linear head baseline, to assess which features the encoder captures in the first place.
To this end, we use the mls-eng-speaker-descriptions train and test set (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024). This dataset contains classification tasks for rate, gender, background noise, and other features (more details in Appendix C).
The results demonstrate that models with contrastive and ASR-based pretraining perform comparably to the plain HuBERT model in predicting paralinguistic annotations, indicating that alignment-focused pretraining does not significantly affect the retention of paralinguistic information. Table 12 shows the accuracies for the different classification tasks for the different models.
@@
Model Speaking Gender Noise Reverberation Speech SDR PESQ
Rate Monotony Noise Speech Quality Avg.
random 14.29 50.00 14.29 20.00 20.00 16.67 16.67 22.25
HuBERT + Class. Head 57.38 64.69 27.89 28.12 39.35 42.85 27.22 41.07
no pretrain 57.07 64.26 27.62 27.38 37.60 33.86 23.72 38.79
ASR pretrain 57.23 57.95 28.87 29.05 36.35 30.19 21.86 37.36
contr-cos-all 56.46 62.17 28.60 29.27 35.39 34.17 24.12 38.60
contr-wasser-all 56.12 62.30 28.26 27.57 37.70 34.25 24.30 38.64
contr-cos-all + asr 55.00 61.29 28.73 26.24 35.29 30.94 24.86 37.48
@@
10% ASR pretrain 12.21 24.822 75.704 49.478 63.358 39.934
   + meta 12.11 24.91 75.702 49.472 63.31 40.154
10% contr-cos-all 13.06 27.188 78.285 60.482 73.9 67.388
   + meta 13.53 27.099 78.213 61.024 74.745 66.932
contr-wasser-all 12.92 29.072 80.522 64.036 77.258 85.017
   + meta 11.97 29.283 80.475 66.124 78.25 88.915
10% contr-cos-all + asr 11.68 27.608 78.94 64.157 76.396 78.577
   + meta 12.01 27.817 78.917 63.112 75.981 76.903
contr-wasser-all + asr 11.23 29.064 80.26 66.727 77.565 88.802
   + meta 11.16 29.342 80.547 65.402 78.26 91.604

### v2

@@ 5.4 Additional pretraining data
As before, we use a 10% finetuning subset to simulate a low-resource scenario. In preliminary experiments with lightweight contrastive pretraining on the embedding layer (Appendix H), we find that increasing the pretraining data leads to much larger improvements for contr-cos loss ($+32.45$ Norm. avg.) than for contr-wasser loss ($+6.60$). Therefore, we report contr-cos results for subsequent experiments; results for contr-wasser can be found in Appendix H. We perform contrastive learning on all layers with and without ASR loss, as these settings yielded the best results in the previous sections. As shown in Table 4, both models contr-cos-all+giga and contr-cos-all+asr+giga achieve an overall improvement of more than $19$ points (Norm. avg.) compared to the same models trained with less pretraining data, reaching $97.12$ and $102.15$ points respectively. Remarkably, despite being finetuned on only 10% of the task-specific data, these models surpass both specialized models, the Seamless model (Communication et al., 2023) for ST and the You et al. (2022) model for SQA. Moreover, they outperform the ASR-pretraining baseline even on the ASR task.
@@
We evaluate our best-performing models in a higher resource setting. We finetune contr-cos-all+giga and contr-cos-all+asr+giga on 100% of the task-specific data. As expected, this leads to even stronger results. But while combining the contrastive and ASR losses provides significant improvements in the low-resource settings, when finetuning with 100% of the available data, the contrastive loss alone performs on par with the combined loss. We hypothesize that the ASR loss is more sensitive to the out-of-domain Giga pretraining data than the contrastive loss, leading to no improvement on the ASR task.
We compare these models against blsp_lsm_7b (Wang et al., 2024a), which has a similar setup to ours in that it only trains the projector while keeping other components frozen, and Qwen2-Audio-7b-Instruct (Chu et al., 2024), a strong baseline SpeechLLM. Notably, BLSP is finetuned on Must-C translation data, whereas for all other tasks, BLSP and Qwen2-Audio results reflect zero-shot performance. Across all tasks, our model outperforms both of these SpeechLLMs.
Results for contr-wasser experiments can be found in Appendix H, these yield similar results but do not outperform contr-cos.
@@ 5.5 Impact of text-speech alignment on capturing paralinguistic features
Speech data is much richer than a sequence of words, and the pretrained speech encoder (HuBERT) might capture additional paralinguistic features. By aligning the richer speech embeddings more closely with text embeddings, the pretraining methods we explore could potentially train the projector to discard these additional features.
Model Speaking Gender Noise Reverberation Speech SDR PESQ
Rate Monotony Noise Speech Quality Avg.
random 14.29 50.00 14.29 20.00 20.00 16.67 16.67 22.25
HuBERT + Class. Head 57.38 64.69 27.89 28.12 39.35 42.85 27.22 41.07
no pretrain 57.07 64.26 27.62 27.38 37.60 33.86 23.72 38.79
ASR pretrain 57.23 57.95 28.87 29.05 36.35 30.19 21.86 37.36
contr-cos-all 56.46 62.17 28.60 29.27 35.39 34.17 24.12 38.60
contr-wasser-all 56.12 62.30 28.26 27.57 37.70 34.25 24.30 38.64
contr-cos-all + asr 55.00 61.29 28.73 26.24 35.29 30.94 24.86 37.48
Model Avg. Gender COMET
Acc in % Correct Diff
HuBERT + LLama 72.21 77.24 0.69
no pretrain 70.61 70.07 0.61
ASR pretrain 74.11 77.46 0.84
contr-cos-all 78.42 82.20 1.09
contr-wasser-all 76.87 81.46 1.01
contr-cos-all + asr 76.10 80.01 0.95
contr-wasser-all + asr 76.21 81.40 0.99
Table 6: Results for the MuST-SHE dataset (Bentivogli et al., 2020) for different models. Following Bentivogli et al. (2020), we evaluate the overall gender accuracy and the COMET score difference between the scores of correctly and incorrectly gendered references. Results are average over three language pairs, individual results can be found in Table 17.
To assess whether this occurs, we examine our contrastively pretrained models on a) a paralinguistic classification task and b) a generation task where paralinugistic information is needed.
For a) we finetune our models on paralinguistic classification tasks before and after pretraining. In addition, we use the speech encoder (HuBERT) with a linear classification head as a baseline. The results demonstrate that models with contrastive and ASR-based pretraining perform comparably to the baseline HuBERT model in predicting paralinguistic annotations, indicating that alignment-focused pretraining does not affect the retention of paralinguistic information significantly. Detailed accuracy results are presented in Table 5. In addition, Table 16 shows that incorporating these features as auxiliary finetuning tasks does not notably alter model performance.
For b) we test our models on the translation dataset MuST-SHE (Bentivogli et al., 2020), where the speakers’s gender determines correct translations. We include a cascaded model (HuBERT + Llama) as a baseline where all paralinguistic features are lost due to the cascaded setup. We find that the gender accuracy of the translations is higher for contrastively pretrained models than for ASR pretrained models or the cascaded model as can be seen in Table 6, further supporting that contrastive pretraining does not hinder a model’s ability to capture paralinguistic features.
@@ 7 Limitations
We analyze whether contrastively pretrained models, designed to align speech and text, retain their ability to capture meta-speech information such as speaker characteristics or noise. While we find that these models do not lose this capacity in both classification and generation tasks, extending this analysis to additional generation tasks would provide a deeper understanding of their impact.
@@
Potential Risks:
Enabling Language Models to understand speech makes them more easily and widely accessible. However, this also lowers the barrier for the potential misuse these models. In addition, SpeechLLMs can reproduce biases seen in the training data. We try to minimize this risk by not training the LLM, so any safeguard measures trained into the LLM also apply to the SpeechLLM. However, the model could still be affected by adversarial attacks.
@@ Acknowledgments
Appendix Overview
The appendix includes the following information:
  • Hyperparameter and model details (§A)
  • Data details (§B)
  • Prompt details (§C)
  • Contrastive training over multiple layers (§D)
  • Details on the mixed-nwp pretraining (§E)
  • Details on combining pretraining losses (§G)
  • Details on pretraining with more data (including translation results for all language pairs) (§H)
  • Results for capturing paralinguistic features (§I) on classification as well as generation tasks
Appendix A Hyperparameters and Model Details
@@
Q-Former Num Attention Heads 12
@@ Appendix B Data
For ASR and ST, we use the MustC-v1 dataset (Di Gangi et al., 2019). We use the en-de portion as English ASR data, and we use language pairs en-de, en-fr, en-it, and en-es for ST. We exclude examples longer than 45 seconds for training and use the tst-common and tst-he test sets for evaluation.
@@
Contrastive pretraining aligns text and speech, which could potentially lead to the loss of paralinguistic features in the embeddings. Therefore, we also train and test our models on mls-eng-speaker-descriptions (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024). We use the mls-eng-speaker-descriptions train and test set. The available training data is quite large, so we subsample it to make training feasible. Since the test set is predefined, we sample the training data such that the test data size is 5% of the data. We let the model predict the following features: Speaking rate, gender, noise, reverberation, speech monotony, SDR noise, and PESQ speech quality. Moreover, we test our models on MuST-SHE (Bentivogli et al., 2020), a translation test set, where the speaker gender is needed to translate the segment correctly. We use the en-es (), en-it and en-fr portion of MuST-SHE.
Details on prompts can be found in Appendix C.
Appendix C Prompts
@@
(a) contr-cos pretrained models
(b) contr-wasser pretrained models
Figure 2: Analysis of contrastive pretraining performed on different layers. We report the normalized average over tasks as in Eq. 3 for contrastive pretraining and finetuning on 10% of task-specific data. The last point on the x-axis refers to contrastive pretraining using the loss summed over all layers.
@@
(a) Loss during contr-cos-petraining
(b) Loss during contr-wasser-petraining
Appendix D Contrastive Loss over Multiple Layers
FT Model ASR ST SQA Contr loss on test Norm.
[…]
N/A Specialized 06.54 30.99 80.02 64.19 77.10 N/A N/A 100
N/A HuBERT + Llama 18.38 19.85 73.92 36.15 54.76 N/A N/A 0
100% no pretrain 12.09 28.84 79.94 64.13 76.53 1.37 1.26 83.08
10% contr-cos-all (multiple of 5) 13.06 27.19 78.29 60.48 73.90 1.08 0.91 67.39
contr-cos-all 13.08 27.25 78.39 61.45 74.91 1.14 1.05 69.41
Table 11: Initial ablations show that calculating the loss over multiples of 5 layers approximates calculating the loss over all layers efficiently. Results of models with contrastive pretraining using cosine similarity (contr-cos) on all (-all) layers, followed by fine-tuning on a 10% subset of task-specific data. The specialized baselines are Whisper (Radford et al., 2022) for ASR, Seamless (Communication et al., 2023) for ST, and You et al. (2022) for SQA. Metrics with * contribute to the normalized average as in Eq. 3.
Contrastive Loss over Different Layers
For the contrastively pretrained model, we experiment with calculating the loss over different layers (Fig. 2), using the final representation of each layer in the LLM. We find that applying contrastive loss to deeper layers significantly increases downstream performance, however, the best layer to use is not consistent. Thus, using a general approach, such as calculating the loss over multiple layers, is the most likely to be robust to different settings. However, calculating the loss over all layers is computationally expensive. Specifically, when using the more computationally expensive Wasserstein distance, the total pretraining time increases by a factor of 4 when applied to all layers. In contrast, for the more efficient cosine similarity, the runtime only increases by 10%. Initial experiments showed that using every 5th layer approximates the results well (Table 11), so we adopted this more efficient method for all experiments.
Contrastive Loss over Different Layers on Different Tasks
Further analysis reveals that contrastive learning on deeper layers benefit the ST and SQA task the most, while improvements for ASR are less consistent. Fig. 3 shows the finetuning performance on 10% on task-specific data after pretraining on different layers for the tasks ASR, ST and SQA.
Contrastive Loss Analysis during Pretraining
Fig. 4 shows the contrastive loss in the pretraining phase. The later the layer, the lower the contrastive loss, with an exception of the embedding layer. For cosine contrastive training, these differences are significantly bigger than for wasserstein contrastive training. Here the losses for pretraining seem to correlated mostly with the performance of the model after finetuning: Models perform better after contrastive pretraining at later layers. However, for ASR, this trend is not so noticeable.
Position Shifts during Pretraining
@@ Appendix D Contrastive Loss over Multiple Layers
10% no pretrain 23.78 19.87 69.72 34.72 46.07 -51.13
ASR pretrain 12.21 24.82 75.70 49.48 63.36 -39.93
10% mixed-nwp w/o punctuation (p) 14.78 23.85 75.78 47.35 66.39 -37.69
mixed-nwp w/o punctuation (k) 15.27 25.22 76.05 53.57 67.68 -39.67
mixed-nwp with punctuation (k) 15.76 24.23 75.09 47.65 61.36 -23.59
mixed nwp w/o punctuation + giga (k) 14.91 26.07 77.17 56.99 70.94 -51.67
10% contr-cos-emb + mixed (k) 14.26 25.96 76.61 51.97 68.40 -46.61
contr-wasser-emb + mixed (k) 14.70 24.15 75.05 48.01 61.48 -26.53
contr-cos-last + mixed (k) 19.02 23.02 73.92 46.29 58.56 -03.86
contr-wasser-last + mixed (k) 16.03 23.09 74.14 43.84 58.20 -12.95
contr-cos-all + mixed (k) 17.63 22.90 73.83 42.15 55.12 -02.14
Table 12: Ablation studies for models with mixed speech and text input. Metrics with * contribute to the normalized average as in Eq. 3. We experiment with two different forced alignment algorithms, one by Pratap et al. (2023, (p)) and one by Kurzinger et al. (2020, (k)). Since the latter performs better in our scenario, we use this aligner for all subsequent experiments.
To address positional mismatches between pretraining and finetuning (detailed in Appendix A), we experiment with position shifts during pretraining. Specifically, we experiment with varying the absolute starting position in RoPE by randomly sampling it during pretraining for each example. However, these do not consistently improve performance (Fig. 2).
Appendix E Mixed Next Word Prediction
For alignment between text and speech, we utilize the forced alignment algorithm described by Kurzinger et al. (2020), which leverages wav2vec 2.0 (Baevski et al., 2020). We also experimented with the alignment algorithm by Pratap et al. (2023), which performed slightly worse in our scenario (Table 12).
@@ Appendix F Mixed Speech-Text Inputs
For the mixed speech-text input, we experiment with two different forced alignment algorithms, one by Pratap et al. (2023, (p)) and one by Kurzinger et al. (2020, (k)). Since the latter performs better in our scenario, we use this aligner for all subsequent experiments.
Pretraining with mixed speech-text next word prediction (nwp) achieves similar results as ASR pretraining. Here, removing punctuation seems to help the model understand the NWP task better.
We also test contrastive pretraining with mixed speech-text input (contr-*+mixed) and find that aligning the representation on the embedding layer works best (however, contrastive learning without mixed input still works better).
@@
10% ASR pretrain 12.21 24.82 75.70 49.48 63.36 39.93
10% contr-cos-emb 13.98 25.47 76.32 54.86 68.49 46.00
   + asr loss 13.18 25.77 76.89 56.12 68.67 51.64
   + mixed + nwp 14.68 25.28 76.18 48.86 63.92 36.41
contr-wasser-emb 15.99 26.10 77.09 54.64 67.83 43.52
   + asr loss 13.19 26.55 77.40 59.02 70.44 57.03
   + mixed + nwp 16.00 23.98 74.59 45.00 59.11 16.87
10% contr-cos-all 13.06 27.19 78.29 60.48 73.90 67.39
   + asr loss 11.68 27.61 78.94 64.16 76.12 78.17
   + mixed + nwp 22.08 23.16 74.20 48.47 62.13 02.07
contr-wasser-all 12.92 29.07 80.52 64.06 77.22 84.96
   + asr loss 11.23 29.06 80.26 66.73 77.19 88.24
   + mixed + nwp 15.72 24.13 75.00 51.21 64.25 27.52
@@ Appendix H Adding more pretraining data
N/A Specialized 06.54 30.99 80.02 64.19 77.10 N/A N/A 100
N/A BLSP-lslm-7b 44.51 28.70 78.68 05.60 21.82 N/A N/A -96.72
@@
   + giga 10.94 29.95 81.21 65.45 79.12 1.06 0.71 097.12
contr-wasser-all 12.92 29.07 80.52 64.06 77.22 1.07 0.64 084.96
   + giga 12.29 29.30 80.79 68.13 80.29 1.09 0.60 092.75
contr-cos-all + asr 11.68 27.61 78.94 63.81 76.12 1.08 0.91 078.17
   + giga 11.12 29.90 81.29 72.16 82.52 0.98 0.65 102.15
contr-wasser-all + asr 11.23 29.06 80.26 66.29 77.19 1.09 0.66 088.24
   + giga 15.13 30.10 81.45 70.47 82.21 1.03 0.30 091.26
100% ASR pretrain 10.28 30.31 80.98 69.87 80.92 1.37 1.22 100.24
   + giga 11.48 30.46 81.03 72.25 82.56 1.37 1.23 099.77
contr-cos-all 12.56 30.69 81.48 71.71 82.50 1.25 1.02 099.06
   + giga 09.31 31.36 81.98 75.12 84.67 1.10 0.83 114.18
contr-wasser-all 10.35 30.62 81.59 68.65 80.21 1.24 0.91 102.46
   + giga 12.10 30.90 81.65 75.49 84.52 1.25 0.87 104.32
contr-cos-all + asr 10.04 30.82 81.56 72.40 82.47 1.19 1.03 106.57
   + giga 10.03 31.54 81.99 76.11 84.94 1.10 0.79 112.66
contr-wasser-all + asr 12.08 30.75 81.62 73.13 83.24 1.33 0.93 102.31
   + giga 09.50 31.20 82.00 74.27 84.39 1.11 0.47 113.37
@@
Results on 100% of the data
The results on 100% of the data can be found in Table 4
Results on different translation directions
[…]
Data en-de en-es en-fr en-it en-de en-es en-fr en-it
N/A Seamless 26.25 33.10 36.49 28.15 79.03 80.38 79.92 80.75
N/A BLSP-lslm-7b 23.30 27.40 31.90 23.20 76.60 78.70 77.70 78.20
N/A Qwen2-Audio-7b 19.60 21.02 27.83 17.83 74.44 74.47 75.69 74.65
10% ASR pretrain 20.31 26.94 28.34 21.71 74.32 76.30 74.56 76.03
   + giga 23.25 29.81 32.42 24.28 77.56 79.12 77.68 79.17
contr-cos-all 23.23 29.20 32.05 24.27 77.42 78.83 77.67 79.22
   + giga 25.29 32.68 35.27 26.55 80.44 81.64 80.97 81.79
contr-cos-all + asr 23.51 29.85 32.62 24.46 78.25 79.42 78.43 79.66
   + giga 25.56 32.34 35.01 26.67 80.63 81.70 80.97 81.87
100% ASR pretrain 25.92 32.61 35.64 27.07 80.20 81.36 80.64 81.70
   + giga 26.05 32.55 35.90 27.33 80.40 81.25 80.61 81.87
contr-cos-all 26.06 33.20 36.09 27.40 80.91 81.88 81.07 82.04
   + giga 26.63 33.81 36.90 28.20 81.23 82.30 81.74 82.63
contr-cos-all + asr 26.12 33.23 35.98 27.93 80.86 82.02 81.24 82.11
   + giga 26.88 33.64 37.15 28.49 81.37 82.34 81.70 82.56
Table 15: Comparison of translation performance of pretraining only on Must-C data or the combination with the bigger Giga dataset (+giga). We compare our models against BLSP (Wang et al., 2024a) and the Qwen2-Audio-7B-Instruct model (Chu et al., 2024).
For the models we discuss in Section 5.4, we also report detailed results for the translation tasks. Table 15 contains the results for the different translation directions.
@@
Speech data is much richer than a sequence of words, and the pretrained speech encoder (HuBERT) might capture additional paralinguistic features. However, contrastive and ASR pretraining could potentially lead to discarding these paralinguistic features, when pulling speech embeddings closer to text embeddings. To assess this, we test our models on a) on paralinguistic classification tasks and b) a generation task where paralinguistic information is needed.
Classification Task on Paralinguistic Features
For a) we finetune our models on paralinguistic classification tasks before and after pretraining. We also compare our models against a HuBERT + linear head baseline, to assess which features the encoder captures in the first place.
To this end, we use the mls-eng-speaker-descriptions train and test set (Pratap et al., 2020; Lacombe et al., 2024; Lyth and King, 2024). This dataset contains classification tasks for rate, gender, background noise, and other features (more details in Appendix B). The results demonstrate that models with contrastive and ASR-based pretraining perform comparably to the plain HuBERT model in predicting paralinguistic annotations, indicating that alignment-focused pretraining does not significantly affect the retention of paralinguistic information. Table 5 shows the accuracies for the different classification tasks for the different models.
@@
Generation Task that Requires Paralinguistic Features
For b) we test our models on MuST-SHE (Bentivogli et al., 2020), a translation dataset where the speaker’s gender is necessary to correctly translate. We test the same models as for a) but also include a cascaded model (HuBERT + Llama31Instruct) as a baseline where all paralinguistic features are lost due to the nature of the cascaded model.
The results can be found in Table 17. Following Bentivogli et al. (2020), we report the average gender accuracy using the evaluation script that comes with the data, and sensitivity, the score difference between the scores of correctly and incorrectly gendered references. We use the COMET difference (COMET Diff) here.
Across all language pairs, the contrastive models consistently outperform the cascaded model and the ASR pretrained model in terms of gender accuracy. On average across the language pairs, ASR pretraining performs 1.9% points above the baseline, whereas the best performing contrastive model (cos-all) performs 6.2% points above the cascaded baseline. The COMET Diffs are also larger for the contrastively pretained models than for ASR pretraining and the baseline.
This underlines our findings that contrastive learning does not mitigate a model’s ability to reason about paralinguistic features. In contrast, ASR pretraining is not as effective at preserving this ability.
@@
10% ASR pretrain 12.21 24.82 75.70 49.48 63.36 39.93
   + meta 12.11 24.91 75.70 49.47 63.31 40.15
10% contr-cos-all 13.06 27.19 78.29 60.48 73.90 67.39
   + meta 13.53 27.10 78.21 61.02 74.75 66.93
contr-wasser-all 12.92 29.07 80.52 64.06 77.22 84.96
   + meta 11.97 29.28 80.48 66.12 78.25 88.92
10% contr-cos-all + asr 11.68 27.61 78.94 63.81 76.12 78.17
   + meta 12.01 27.82 78.92 63.11 75.98 76.90
contr-wasser-all + asr 11.23 29.06 80.26 66.29 77.19 88.24
   + meta 11.16 29.34 80.55 65.40 78.26 91.60
@@
Lang. Pair Model Avg. Gender Acc in % COMET correct COMET diff
en-es HuBERT + LLama 73.52 78.69 0.7882
no pretrain 69.67 70.89 0.6510
ASR pretrain 73.19 78.27 0.8523
contr-cos-all 77.64 82.97 1.1382
contr-wasser-all 76.73 82.13 1.0784
contr-cos-all + asr 76.14 80.67 1.0183
contr-wasser-all + asr 75.50 82.30 1.0809
en-fr HuBERT + LLama 70.15 75.13 0.5278
no pretrain 70.34 68.66 0.4591
ASR pretrain 72.64 75.94 0.6852
contr-cos-all 77.83 82.92 1.1118
contr-wasser-all 75.48 79.97 0.8856
contr-cos-all + asr 75.73 78.57 0.8594
contr-wasser-all + asr 75.25 79.96 0.8909
en-it HuBERT + LLama 72.96 77.90 0.7401
no pretrain 71.83 70.66 0.7281
ASR pretrain 76.50 78.16 0.9684
contr-cos-all 79.79 80.71 1.0103
contr-wasser-all 78.40 82.29 1.0551
contr-cos-all + asr 76.44 80.78 0.9682
contr-wasser-all + asr 77.90 81.93 0.9858
Table 17: Results for the MuST-SHE dataset (Bentivogli et al., 2020) for different models. Following Bentivogli et al. (2020), we evaluate the overall gender accuracy (using the evaluation script available with the dataset), and the score difference between the scores of correctly and incorrectly gendered references. We use the COMET difference (COMET Diff) here.